The tub of a washer goes into its spin-dry cycle, starting from rest and reaching an angular speed of 6.0 rev/s in 9.0 s. At this point, the person doing the laundry opens the lid, and a safety switch turns off the washer. The tub slows to rest in 10.0 s. Through how many revolutions does the tub turn during this 19-s interval? Assume constant angular acceleration while it is starting and stopping.
Answer #1

Average angular speed in the first 9 s

= (0 + 6)/2 = 3 rev/s

Therefore, no. of revolutions in the first 9 s = 3* 9 = 27

Average angular speed in the next 10 s

= (6 + 0)/2 = 3 rev/s

Therefore, no. of revolutions in the next 10 s = 3 * 10 = 30

Therefore total no. of revolutions in the 19 s = 27 + 30 = 57